OpenStudy (anonymous):

ok this is a long one I will have to do it in 3 parts Use x to represent trucks with automatic transmissions and y to represent those with standard transmissions. Write an equation that expresses this situation. I got x+y=136 then As the sales person takes you around the lot, he mentions that there are 3 times as many trucks with automatic transmissions than with standard. Using the same variables as in question 1, write an equation that expresses this fact. I got 3x+y=136 now how do i get Use the equations from questions 1 and 2 to write a system of equations. Use the substitution method

OpenStudy (amistre64):

ow ow ow ow ow....my eyess!!!, can you clean this up any :) Im sure there are 2 equations in it that can be "systemized".

OpenStudy (amistre64):

id use A and S as the variables really...

OpenStudy (amistre64):

how can x+y = 136 that would mean that the total number of trucks is 136. then you get 3(x) + y = 136......it doesnt make sense. you increased the amount of x by a factor of "3" but still have the same number of trucks... the only way that is possible is if there are no "x" trucks

OpenStudy (anonymous):

the problem states there are 146 cars total

OpenStudy (amistre64):

then x+y = 146 in total? right?

OpenStudy (anonymous):

i am so sorry 136

OpenStudy (amistre64):

.......i can work it with either number, i just need to know which one to use :)

OpenStudy (anonymous):

x+y=136 3x+y=136 this needs to be done in subtraction method

OpenStudy (amistre64):

Ima use A and S....for auto and standard.... ok? A + S = 136 A = 3S this is the info we got

OpenStudy (amistre64):

there are 3 times as many automatics as standards....so A = 3S..right?

OpenStudy (amistre64):

3S + S = 136 4S = 136 S = 136/4 S = 34 A = 102

OpenStudy (anonymous):

Yes there are 3 times as many automatic than standard trucks. I dont know if i worte it out correct. I had 3A+S=136 but i dont think this seems to be right

OpenStudy (amistre64):

it aint right :) total trucks is 136...you cant have 3 times as many MORE trucks and still have 136 unless there were NO trucks of that kind to begin with.... to I assume there is an error

OpenStudy (amistre64):

if anything: A - 3S = 0

OpenStudy (anonymous):

ok thank you so much. I have one more question for you how did you get A=102 did you just subtract the 136- 34?

OpenStudy (amistre64):

yup....